# kabr-tools: Automated framework for multi-species behavioral monitoring

### v1

@@ 1 Introduction
Recent advances in remote sensing and computational analysis offer transformative opportunities for scaling behavioral monitoring to meet ecological research needs (Hughey et al., 2018; Corcoran et al., 2021; Besson et al., 2022; Koger et al., 2023; Kline et al., 2025b). Drone technology provides aerial observation platforms that can simultaneously monitor multiple individuals across groups while minimizing observer effects and visual obstructions (Afridi et al., 2024; Pedrazzi et al., 2025; Schad and Fischer, 2023). More critically, the integration of computer vision and machine learning enables automated processing of the large video datasets generated by drone surveys, transforming behavioral ecology from observation-limited to data-intensive science (Petso et al., 2021; Zhang et al., 2024; Saoud et al., 2024; Axford et al., 2024; Chan et al., 2024; Elhorst et al., 2025; Kline et al., 2025a). These computational approaches can process behavioral information at scales previously impossible with manual methods, enabling more efficient landscape-level monitoring, multi-species comparative studies, and long-term assessment of behavioral responses to environmental change.
@@
We demonstrate the effectiveness of drones in collecting data, paired with machine learning for analysis, by benchmarking the KABR dataset against traditional, ground-based expert observation methods. We validate our expanded framework using five key metrics and three case studies drawn from the behavioral data in the original KABR video clips and 50 minutes of newly released full-length videos. We provide worked examples for fourteen new videos, including annotated detections, tracks, behavior labels, and telemetry data. Our enhanced framework provides automated extraction of time budgets, behavioral transitions, social interactions, and habitat associations across these diverse study systems. The modular design enables integration of emerging machine learning models and adaptation to new species and study contexts. We demonstrate the ecological applications of this approach through three case studies, which include the computational analysis of anti-predator behavior in Grevy’s zebras, transitions in zebra behavior, and inter- and intra-specific social interactions. Our analysis reveals species-specific risk management strategies across habitat gradients. We provide comprehensive documentation  [^1] https://imageomics.github.io/kabr-tools/, open-source code  [^2] https://github.com/Imageomics/kabr-tools. All the data is available on HuggingFace, including methodological comparisons [^3] https://huggingface.co/datasets/imageomics/kabr-methodology, and worked examples for three sessions, including full length videos full-length videos [^4] https://huggingface.co/datasets/imageomics/kabr-full-video, detections, tracks, behavior annotations, and associated telemetry [^5] https://huggingface.co/datasets/imageomics/kabr-worked-examples. We provide a glossary bridging behavioral ecology and computer vision terminology to facilitate interdisciplinary adoption (Tab. 1). This evolution from dataset creation to a comprehensive computational ecology framework advances automated wildlife monitoring capabilities while providing scalable tools for addressing contemporary conservation biology and ecosystem management challenges.
@@ 2.3.2 Ecological Tasks and Case Studies
We focus on five broad ecological tasks: time budget analysis, behavioral transitions, social interactions, land use and movement patterns, and social group composition analysis, summarized in Tab. 4, (Hughey et al., 2018; Smith and Pinter-Wollman, 2021; Pedrazzi et al., 2025). These tasks provide a multifaceted understanding of animal behavior by capturing not just how much time animals spend in different activities, but also how behaviors transition over time, vary across habitats, and are shaped by social context. Time budgets and transition matrices offer standardized metrics to quantify behavioral patterns and flexibility, while drone-enabled detection of interactions and movement pathways reveals dynamics often missed by ground observers. Together, these methods help ecologists link behavior to environmental and social drivers, which is essential for interpreting animal responses to habitat change, human disturbance, and conservation interventions.
@@ 3.1 Comparison of Behavior Data Collection Methods
To evaluate the effectiveness of drone-based behavioral sampling, we perform three key comparisons: (1) ground-based scan versus ground-based focal sampling (Section 3.1.1), (2) ground-based focal versus drone-based focal sampling (Section 3.1.2), and (3) manual versus automated behavior annotations of drone-based observations (Section 3.1.3). We use confusion matrices to compare concordance between these methodologies (Fig. 6). Confusion matrices are tabular summaries that quantify agreement between two classification methods by reporting the frequency with which each behavior label from one method matches or differs from the corresponding label in the other. Our comparison confirmed that scan sampling captures concurrent behavior of the entire herd, but captures fewer behavior categories overall compared to individual focal sampling. Drone-based focal sampling reduced out-of-sight time by 14.7% and captured more behavioral transitions compared to ground-based focal sampling. A comparison of manual versus automated drone annotations showed high agreement for dominant behaviors, confirming the utility of automated pipelines while revealing systematic errors in rare or subtle behaviors. Together, these comparisons demonstrate that drone-based sampling with machine learning provides a powerful complement to traditional methods, combining the comprehensive coverage of scan sampling with the fine granularity of focal sampling, while enabling scalable, automated behavioral analysis at high resolution.
@@ 3.2.1 Case Study 1: Grevy’s Zebras Landscape of Fear
We analyze anti-predator strategies in Grevy’s zebras using 83 mini-scenes. We selected mini-scenes from six representative sessions, containing a mix of herd sizes and habitat types (Tab. 10). To ensure sufficiently long behavior sequences for analysis, we selected mini-scenes $\geq 60$ seconds. We classify the mini-scenes by habitat type based on vegetation density by visually examining the mini-scene footage to classify habitat type as open (little to no vegetation) or closed (dense vegetation). Examples of habitat categories are illustrated in Fig. 8. The median herd size is three individuals; herds $\leq 3$ are categorized as small, herds >3>3 are large. For each mini-scene, the time budget was calculated as a proportion of Walk, Head Up, Graze, and Other. We use Head Up behaviors to indicate instances of vigilance, where animals survey their habitat for potential risks.
@@
(a) Example of closed habitat with dense vegetation (Session 4 from Tab. 10)
(b) Example of open habitat with with little vegetation (Session 5 from Tab. 10)
@@ 3.2.2 Case Study 2: Zebra State Shifts
We analyzed 748 mini-scenes, totaling 2 hours 30 minutes, containing zebra locomotion behaviors of Graze, Walk, Trot, and Run. We quantify the likelihood of zebras transitioning between behaviors (Fig. 10) by normalizing the transitions by the count of occurrences, down-sampled every 10 seconds. Individuals are most likely to remain in the same behavioral state, with high self-transition probabilities for Graze (91.1%), Walk (83.0%), and Run (68.4%). Previous instances of Trot were most likely to transition to Walk (53.4%). Transitions from relaxed states to more alert behaviors are also evident: for instance, individuals in a Walk state transition to Trot 2.6% of the time and to Run 2.0% of the time, while Trot transitions to Run with a relatively high probability of 14.3%. These transition dynamics provide insight into how animals escalate movement in response to internal or external stimuli and offer a foundation for modeling behavioral state changes. Understanding these patterns is critical for developing behavior-adaptive autonomous drones that can anticipate transitions, such as from grazing to fleeing, and adjust drone flight paths accordingly to minimize disturbance while maximizing data capture.
@@ 3.2.3 Case Study 3: Cross-Species Encounters
A mixed species herd of Grevy’s zebras, Plains zebras, and giraffes was observed grazing together (Fig. 11). From the drone videos, 157 mini-scenes were generated and automatically labeled with behaviors using the KABR X3D model (Kholiavchenko et al., 2024c). To assess patterns of spatial proximity, we computed normalized bounding box overlap counts between individuals of the same and different species (Tab. 7). Normalization was performed by dividing the total number of overlaps by the number of possible pairs for each species combination. Both zebra species showed higher overlap compared to giraffes. Grevy’s zebras exhibited a high normalized overlap rate (87.93 overlaps per pair), suggesting frequent and sustained close proximity among group members. Plains zebras showed the highest normalized value (93.00), though this reflects a single dyad and should be interpreted with caution due to the small sample size. Giraffes had a moderate normalized overlap rate (26.00), consistent with a more spatially dispersed grouping structure. Cross-species overlaps were rare. The normalized overlap between Grevy’s and plains zebras was low (1.27), and no overlaps were observed between giraffes and zebras (0.00), indicating strong spatial segregation between species in the observed dataset. Overall, these patterns suggest that intra-species proximity is most prominent in Grevy’s zebras, with limited interaction between species.
@@
Figure 11: Image from mini-scene of mixed species herd consisting of $1$ Grevy’s zebras (center), 2 Plains zebras (left), and 3 reticulated giraffes (left, center, right) analyzed for Case Study 3: Cross-Species Encounters.
@@ Acknowledgments
This work is supported by the National Science Foundation under Award No. 2118240 "HDR Institute: Imageomics: A New Frontier of Biological Information Powered by Knowledge-Guided Machine Learning" and Award No. 2112606 "Intelligent Cyberinfrastructure with Computational Learning in the Environment (ICICLE)".
@@ 6 Data Availability Statement
All code, annotated datasets, and documentation are publicly available at https://github.com/Imageomics/kabr-tools. Documentation in Github includes annotation guidelines and best practices for managing large-scale ecological video datasets. The data used for analysis is available on HuggingFace and includes full-length videos, worked examples, and methodology comparison.
@@ D Plains male zebra, travelling with plains harem 72 sec
Table 9: Demographic details of individual zebras’ behaviors analyzed to compare ground focal with drone focal annotations. The temporal overlap quantifies the time when ground and drone focal sampling coincide.
@@ 6 Open Large (5) 20/01/23
Table 10: Summary of data used for Case Study 1: Grevy’s Zebras Landscape of Fear, by habitat and herd size. Herds larger than the median (4.2) are considered ‘Large’, all others are considered ‘Small’. For session 3, the habitat type transitioned from Open to Closed as the animals moved to a more dense habitat during the video duration.

### v2

@@ 1 Introduction
Recent advances in remote sensing and computational analysis offer transformative opportunities for scaling behavioral monitoring to meet ecological research needs (Hughey et al., 2018; Corcoran et al., 2021; Besson et al., 2022; Koger et al., 2023; Kline et al., 2025d). Drone technology provides aerial observation platforms that can simultaneously monitor multiple individuals across groups while minimizing observer effects and visual obstructions (Afridi et al., 2024; Pedrazzi et al., 2025; Schad and Fischer, 2023). More critically, the integration of computer vision and machine learning enables automated processing of the large video datasets generated by drone surveys, transforming behavioral ecology from observation-limited to data-intensive science (Petso et al., 2021; Zhang et al., 2024; Saoud et al., 2024; Axford et al., 2024; Chan et al., 2024; Elhorst et al., 2025; Kline et al., 2025a). These computational approaches can process behavioral information at scales previously impossible with manual methods, enabling more efficient landscape-level monitoring, multi-species comparative studies, and long-term assessment of behavioral responses to environmental change.
@@
We demonstrate the effectiveness of drones in collecting data, paired with machine learning for analysis, by benchmarking the KABR dataset against traditional, ground-based expert observation methods (Kline et al., 2025b). We validate our expanded framework using five key metrics and three case studies drawn from the behavioral data in the original KABR video clips and 50 minutes of newly released full-length videos (Kline et al., 2025c). We provide worked examples for fourteen new videos, including annotated detections, tracks, behavior labels, and telemetry data (Kline et al., 2025e). Our enhanced framework provides automated extraction of time budgets, behavioral transitions, social interactions, and habitat associations across these diverse study systems. The modular design enables integration of emerging machine learning models and adaptation to new species and study contexts. We demonstrate the ecological applications of this approach through three case studies, which include the computational analysis of anti-predator behavior in Grevy’s zebras, transitions in zebra behavior, and inter- and intra-specific social interactions. Our analysis reveals species-specific risk management strategies across habitat gradients. We provide comprehensive documentation  [^1] https://imageomics.github.io/kabr-tools/, open-source code  [^2] https://github.com/Imageomics/kabr-tools. All the data is available on HuggingFace, including methodological comparisons [^3] https://huggingface.co/datasets/imageomics/kabr-methodology, and worked examples for three sessions, including full length videos full-length videos [^4] https://huggingface.co/datasets/imageomics/kabr-full-video, detections, tracks, behavior annotations, and associated telemetry [^5] https://huggingface.co/datasets/imageomics/kabr-worked-examples. We provide a glossary bridging behavioral ecology and computer vision terminology to facilitate interdisciplinary adoption (Tab. 1). This evolution from dataset creation to a comprehensive computational ecology framework advances automated wildlife monitoring capabilities while providing scalable tools for addressing contemporary conservation biology and ecosystem management challenges.
@@ 2.3.2 Ecological Tasks and Case Studies
We focus on five broad ecological tasks: time budget analysis, behavioral transitions, social interactions, land use and movement patterns, and social group composition analysis (Hughey et al., 2018; Smith and Pinter-Wollman, 2021; Pedrazzi et al., 2025), summarized in Tab. 4. These tasks provide a multifaceted understanding of animal behavior by capturing not just how much time animals spend in different activities, but also how behaviors transition over time, vary across habitats, and are shaped by social context. Time budgets and transition matrices offer standardized metrics to quantify behavioral patterns and flexibility, while drone-enabled detection of interactions and movement pathways reveals dynamics often missed by ground observers. Together, these methods help ecologists link behavior to environmental and social drivers, which is essential for interpreting animal responses to habitat change, human disturbance, and conservation interventions.
@@ 3.1 Comparison of Behavior Data Collection Methods
To evaluate the effectiveness of drone-based behavioral sampling, we perform three key comparisons: (1) ground-based scan versus ground-based focal sampling (Section 3.1.1), (2) ground-based focal versus drone-based focal sampling (Section 3.1.2), and (3) manual versus automated behavior annotations of drone-based observations (Section 3.1.3) (Kline et al., 2025b). We use confusion matrices to compare concordance between these methodologies (Fig. 6). Confusion matrices are tabular summaries that quantify agreement between two classification methods by reporting the frequency with which each behavior label from one method matches or differs from the corresponding label in the other. Our comparison confirmed that scan sampling captures concurrent behavior of the entire herd, but captures fewer behavior categories overall compared to individual focal sampling. Drone-based focal sampling reduced out-of-sight time by 14.7% and captured more behavioral transitions compared to ground-based focal sampling. A comparison of manual versus automated drone annotations showed high agreement for dominant behaviors, confirming the utility of automated pipelines while revealing systematic errors in rare or subtle behaviors. Together, these comparisons demonstrate that drone-based sampling with machine learning provides a powerful complement to traditional methods, combining the comprehensive coverage of scan sampling with the fine granularity of focal sampling, while enabling scalable, automated behavioral analysis at high resolution.
@@ 3.2.1 Case Study 1: Grevy’s Zebras Landscape of Fear
We analyze anti-predator strategies in Grevy’s zebras using 83 mini-scenes (Kline et al., 2024b, 2025e). We selected mini-scenes from six representative sessions, containing a mix of herd sizes and habitat types (Tab. 10). To ensure sufficiently long behavior sequences for analysis, we selected mini-scenes $\geq 60$ seconds. We classify the mini-scenes by habitat type based on vegetation density by visually examining the mini-scene footage to classify habitat type as open (little to no vegetation) or closed (dense vegetation). Examples of habitat categories are illustrated in Fig. 8. The median herd size is three individuals; herds $\leq 3$ are categorized as small, herds >3>3 are large. For each mini-scene, the time budget was calculated as a proportion of Walk, Head Up, Graze, and Other. We use Head Up behaviors to indicate instances of vigilance, where animals survey their habitat for potential risks.
@@
(a) Example of closed habitat with dense vegetation (Session 4 from Tab. 10), (Kline et al., 2024a).
(b) Example of open habitat with with little vegetation (Session 5 from Tab. 10), (Kline et al., 2025e).
@@ 3.2.2 Case Study 2: Zebra State Shifts
We analyzed 748 mini-scenes (Kline et al., 2024b). The mini-scenes total 2 hours 30 minutes, containing zebra locomotion behaviors of Graze, Walk, Trot, and Run. We quantify the likelihood of zebras transitioning between behaviors (Fig. 10) by normalizing the transitions by the count of occurrences, down-sampled every 10 seconds. Individuals are most likely to remain in the same behavioral state, with high self-transition probabilities for Graze (91.1%), Walk (83.0%), and Run (68.4%). Previous instances of Trot were most likely to transition to Walk (53.4%). Transitions from relaxed states to more alert behaviors are also evident: for instance, individuals in a Walk state transition to Trot 2.6% of the time and to Run 2.0% of the time, while Trot transitions to Run with a relatively high probability of 14.3%. These transition dynamics provide insight into how animals escalate movement in response to internal or external stimuli and offer a foundation for modeling behavioral state changes. Understanding these patterns is critical for developing behavior-adaptive autonomous drones that can anticipate transitions, such as from grazing to fleeing, and adjust drone flight paths accordingly to minimize disturbance while maximizing data capture.
@@ 3.2.3 Case Study 3: Cross-Species Encounters
A mixed species herd of Grevy’s zebras, Plains zebras, and giraffes was observed grazing together, illustrated in Fig. 11) (Kline et al., 2025e). From the drone videos, 157 mini-scenes were generated and automatically labeled with behaviors using the KABR X3D model (Kholiavchenko et al., 2024c). To assess patterns of spatial proximity, we computed normalized bounding box overlap counts between individuals of the same and different species (Tab. 7). Normalization was performed by dividing the total number of overlaps by the number of possible pairs for each species combination. Both zebra species showed higher overlap compared to giraffes. Grevy’s zebras exhibited a high normalized overlap rate (87.93 overlaps per pair), suggesting frequent and sustained close proximity among group members. Plains zebras showed the highest normalized value (93.00), though this reflects a single dyad and should be interpreted with caution due to the small sample size. Giraffes had a moderate normalized overlap rate (26.00), consistent with a more spatially dispersed grouping structure. Cross-species overlaps were rare. The normalized overlap between Grevy’s and plains zebras was low (1.27), and no overlaps were observed between giraffes and zebras (0.00), indicating strong spatial segregation between species in the observed dataset. Overall, these patterns suggest that intra-species proximity is most prominent in Grevy’s zebras, with limited interaction between species.
@@
Figure 11: Image from mini-scene of mixed species herd consisting of $1$ Grevy’s zebras (center), 2 Plains zebras (left), and 3 reticulated giraffes (left, center, right) analyzed for Case Study 3: Cross-Species Encounters (Kline et al., 2025e).
@@ Acknowledgments
This work was supported by the Imageomics Institute, which is funded by the US National Science Foundation’s Harnessing the Data Revolution (HDR) program under Award 2118240 (Imageomics: A New Frontier of Biological Information Powered by Knowledge-Guided Machine Learning). Additional support was provided by the AI Institute for Intelligent Cyberinfrastructure with Computational Learning in the Environment (ICICLE), funded by the US National Science Foundation under Award 2112606. We thank the Mpala Research Centre for their support in collecting the dataset.
@@ 6 Data Availability Statement
All code, annotated datasets, and documentation are publicly available at https://github.com/Imageomics/kabr-tools. Documentation in Github includes annotation guidelines and best practices for managing large-scale ecological video datasets. The data used for analysis is available on HuggingFace and includes full-length videos (Kline et al., 2025c), worked examples (Kline et al., 2025e), and methodology comparison (Kline et al., 2025b).
The data was collected at the Mpala Research Centre in Kenya, in accordance with Research License No. NACOSTI/P/22/18214. The data collection protocol adhered strictly to the guidelines set forth by the Institutional Animal Care and Use Committee under permission No. IACUC 1835F.
@@ D Plains male zebra, travelling with plains harem 72 sec
Table 9: Demographic details of individual zebras’ behaviors analyzed to compare ground focal with drone focal annotations, Kline et al. [2025b]. The temporal overlap quantifies the time when ground and drone focal sampling coincide.
@@ 6 Open Large (5) 20/01/23
Table 10: Summary of data used for Case Study 1: Grevy’s Zebras Landscape of Fear, by habitat and herd size, [Kline et al., 2024a, 2025e]. Herds larger than the median (4.2) are considered ‘Large’, all others are considered ‘Small’. For session 3, the habitat type transitioned from Open to Closed as the animals moved to a more dense habitat during the video duration.